OpenStudy (anonymous):

Can someone verify if I got the correct answer? Thank you √63a^9b * √27a^10b^8 = 9a^9b^3 .√21ab

OpenStudy (anonymous):

\[9a^9b^3\times \sqrt{21ab}\]

OpenStudy (anonymous):

that the answer?

OpenStudy (mertsj):

\[\sqrt{63a ^{9}b}\sqrt{27a ^{10}b ^{8}}\]

OpenStudy (mertsj):

Do I have the original problem right?

OpenStudy (anonymous):

yes..

OpenStudy (mertsj):

Then the answer is:\[9a ^{9}b ^{4}\sqrt{21ab}\]

OpenStudy (anonymous):

Thanks for your help.

OpenStudy (mertsj):

yw